# Avoiding Catastrophe in Continuous Spaces by Asking for Help

### v2

@@ 4.1 Intuition behind the algorithm
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq\frac{6L}{T^{\frac{0.5}{n+0.5}}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
Corollary 4.6 makes the near-zero regret and sublinear mentor queries more explicit.
Corollary 4.6.
Assume $\pi^{*}\in\Pi$, where $\Pi$ has finite Natarajan dimension $d$ and is $\varepsilon(T)^{1/n}$-local for $\varepsilon(T)=T^{\frac{-n}{n+0.5}}$. Then Algorithm 1 with $g(T)=2cdT^{\frac{n}{n+0.5}}\log(d\,T)$ and $r=\varepsilon(T)^{1/n}$ makes at most $O\left(T^{\frac{n}{n+0.5}}(d\log(d\,T)+\operatorname*{vol}(\operatorname*{conv}(S)))\right)$ queries and satisfies
$\lim_{T\to\infty}\>\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]=0$
When $d$ is constant, the number of queries in Corollary 4.6 is $o(T)$. (Actually, $d$ can be as large as $T^{\frac{0.5-\delta}{n+0.5}}$ for $\delta\in(0,1/2)$.) Also, since the expected regret goes to zero, the regret must go to zero with high probability (since regret is always nonnegative).
Finally, note that standard PAC learning lower bounds do not directly yield lower bounds for our setting: even $\Pr[a_{t}\neq\pi^{*}(s_{t})]=1$ is insufficient for high regret, since the payoffs could be nearly optimal. Also, PAC learning lower bounds typically state that there exists a distribution for which the algorithm performs poorly, while our lower bound uses a fixed known distribution.
5 Near-zero regret for simple mentor policies in 1D with adversarial states
Our final result is another algorithm with near-zero regret and sublinear queries, but under different assumptions. In this section, we assume a 1D state space and we allow the state sequence to be adversarially chosen. We consider on the following notion of simplicity:
Definition 5.1.
Given a mentor policy $\pi^{*}$, partition the state space $S$ into intervals such that all states within each interval share the same mentor action. Let $\{S_{1},\dots,S_{k}\}$ be a partition that minimizes the number of intervals. We call each $S_{j}$ a segment. Let $f(\pi^{*})$ denote the number of segments in $\pi^{*}$.
This is similar conceptually to Natarajan dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., optimal actions) on $\mathbf{s}$. The number of segments also relates to the minimum depth needed by a neural network to learn the policy [7]. Figure 1 shows why the Lipschitz constant does not bound the number of segments ($f(T)$ can be arbtrarily large).
We prove the following result:
Theorem 5.2.
For any $c\in[0,1]$, any $\mathbf{s}\in S^{T}$, and any $\pi^{*}$ with $f(\pi^{*})\leq K$, Algorithm 2 with $g(T)=T^{c}$ makes at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries and satisfies $R_{T}(\mathbf{s},\mu,\pi^{*})\leq 6LKT^{1-2c}$.
Corollary 5.3 makes the sublinear queries and near-zero regret of Theorem 5.2 more explicit when the number of segments is a constant.
Corollary 5.3.
For any $c\in(1/2,1)$, Algorithm 2 with $g(T)=T^{c}$ makes $O(T^{c}\operatorname*{len}(\operatorname*{conv}(S)))$ queries and satisfies
$\lim_{T\to\infty}\ \sup_{\mathbf{s}\in S^{T}}\ \sup_{\mu}\ \sup_{\pi^{*}:f(\pi^{*})\leq K}\ R_{T}(\mathbf{s},\mu,\pi^{*})=0$
Our algorithm does not need to know the Lipschitz constant $L$ or the number of segments; it only needs to know the time horizon $T$.
5.1 Intuition behind the algorithm
The algorithm maintains a set of buckets which partition the observed portion of the state space. Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the state space. As long as the bucket contains a query from a prior time step, the Lipschitz constraint allows us to bound $\mu^{*}(s_{t})-\mu(s_{t},a_{t})$ based on the length of the bucket containing $s_{t}$. We always query if the bucket does not contain a prior query
Regardless of the bucket length, the adversary can still place multiple segments in the same bucket $B$. A single query only tells us the optimal action for one of those segments, so we risk a payoff as bad as $\mu^{*}(s_{t})-O(\operatorname*{len}(B))$ whenever we choose not to query. We can endure a limited number of such payoffs, but if we never query again in that bucket, we may suffer $\Theta(T)$ such payoffs. Letting $\mu^{*}(s_{t})=1$ for simplicity, that would lead to $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{1}{O(g(T))}\big{)}^{\Theta(T)}$, which converges to 0 (i.e., guaranteed catastrophe) when $g(T)\in o(T)$.
This approach is reminiscent of Algorithm 1 in that it collects an initial dataset (one query per initial bucket), and then performs supplemental queries as needed based on the observed states. We do not think this is a coincidence.
3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{*}(s)$ for each $s\in S_{Q}$
5:     for $t$ from $1$ to $T$ do
6:         EvaluateState$(s_{t})$
7:function EvaluateState($s\in S$)
10:         $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
12:         EvaluateState$(s)$
13:     else
14:         $B\leftarrow$ any bucket containing $s$
16:              Query mentor and observe $\pi^{*}(s)$
17:              $\pi(s)\leftarrow\pi^{*}(s)$
18:              $S_{Q}\leftarrow S_{Q}\cup\{s\}$
19:              $n_{B}\leftarrow n_{B}+1$
21:              Let $s^{\prime}\in S_{Q}\cap B$
22:              Take action $\pi(s^{\prime})$
23:              $n_{B}\leftarrow n_{B}+1$
24:         $\triangleright$ Bucket is full: split bucket and try again
25:              $B=[x,y]$
27:              $(s_{B_{1}},s_{B_{2}})\leftarrow(0,0)$
28:              $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
29:              EvaluateState$(s)$
Algorithm 2 achieves near-zero regret when the mentor’s policy has a bounded number of segments.
6 Limitations
Our paper contains both positive and negative results. As such, our claim in this paper is not “avoiding catastrophe is possible (or impossible) under assumptions X, Y, Z, and we argue that X, Y, Z hold”. Instead, we aim to understand the conditions under which it is possible to avoid catastrophe. Our claim is “avoiding catastrophe is possible under X, Y, Z, but impossible under A, B, C; we might sometimes expect X, Y, Z and sometimes expect A, B, C”. For example, some real-world instances might exhibit bounded Natarajan dimension, while others might not. Our paper elucidates what to expect from either scenario.
The goal, then, is to make the gap between X, Y, Z and A, B, C as small as possible: to isolate the boundary of tractability. Our results suggest that the key factor determining tractability is the complexity of the mentor policy class. Our negative result (Theorem 3.1) shows that a simple environment (bounded 1D state space, known uniform distribution over states, two actions) is not enough for tractability if the mentor policy class is complex. On the other hand, our primary positive result (Theorem 4.5) holds for quite general environments (unbounded $n$-dimensional state space, unknown arbitrary distribution over states, many actions), as long as the mentor policy class is simple.
One remaining gap between those two results is the assumption of $r$-locality in Theorem 4.5. Although we think $r$-locality is a fairly plausible assumption (as argued in Section 4), that assumption would need to be removed in order for Theorem 3.1 and Theorem 4.5 to fully match.
Another question is how far this positive result generalizes. In particular, the assumptions of (1) only caring about catastrophe and (2) i.i.d. states significantly limit the real-world applicability of our algorithm. In future work, we hope to extend this result to general MDPs.
Our secondary positive result (Theorem 5.2) also suggests that mentor policy complexity could be the key factor determining tractability. However, it has the drawback that the number of segments in the mentor policy is a less established measure of complexity than the well-studied notion of Natarajan dimension. In our opinion, the more serious limitation is the restriction to a 1D state space. Although Theorem 5.2 does permit adversarial state sequences, we feel that Theorem 4.5 does a better overall job of isolating the boundary of tractability.
@@ 7 Conclusion and future work
In this paper, we proposed a model of avoiding catastrophe in sequential decision-making. We showed that in general, any algorithm either constantly asks for help or is guaranteed to cause catastrophe. However, the problem becomes tractable when the mentor’s policy is relatively simple, for two different definitions of “simple”.
There remain some technical questions within this paper’s model: (1) Can Theorem 4.5 be generalized to adversarial state sequences? (2) Is $r$-locality necessary? (3) Are there are reasonable alternatives to the Lipschitz assumption? (Indeed, the utility of $r$-locality steps from the Lipschitz assumption.)
As mentioned, we are also excited to move beyond the current model to full MDPs: recall the open problem we posed in Section 1. Most applications are better captured by MDPs than by this paper’s assumptions of i.i.d. states and only caring about catastrophe.
Finally, although we intentionally avoided Bayesian inference in this paper, it seems premature to abandon that tool entirely. Are there tractable approximations to Bayesian approaches?
@@ Acknowledgements
This work was supported in part by a gift from Open Philanthropy to the Center for Human-Compatible AI (CHAI) at UC Berkeley. This paper also benefited from discussions with many researchers at CHAI. We would like to particularly thank Michael Cohen, Bhaskar Mishra, Cam Allen, and Li Dayan for helpful feedback.
@@ A.1 Proof roadmap
Recall that $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\mathbf{a}}$ denotes the expectation over both the random states and any additional randomness in the algorithm.
@@
    To complete the proof, we observe that $\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu_{f,\mathbf{x}},\pi^{*})$. This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\mathbf{x}}$ has high expected regret, then the worst case $\mu$ also has high expected regret.
@@ A.2 Proof
Independently sample $\mathbf{x}\sim U(\{0,1\}^{f(T)})$ and $s\sim U(S^{T})$. [^13] That is, the entire set $\{x_{1},\dots,x_{f(T)},s_{1},\dots,s_{T}\}$ is mutually independent. Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$,
@@
Part 3: $V_{j}^{\prime}$ is not too small. Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in V_{j}$, so $a_{t}\in\{0,1\}$ for all $t\in V_{j}$. Since the agent chooses one of two actions for each $t\in V_{j}$, the more frequent action must be chosen chosen at least half of the time: $a_{t}=y_{j}$ for at least half of the time steps in $V_{j}$. Since $x_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $a_{t}=y_{j}\neq x_{j}$ for those time steps, so $|V_{j}^{\prime}|\geq|V_{j}|/2$.
@@
$\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ \prod_{j\in\bar{J}}\prod_{t\in V_{j}^{\prime}}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
$\leq$ $\ \prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}^{\prime}|}$
$\leq$ $\ \prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}|/2}$
@@
Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$ for brevity. See 3.1
@@
If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{*}$. Let $f(T)=\sqrt{(g(T)+1)T}$. Then by Lemma A.1, $g(T)\in o(f(T))$ and $f(T)\in o(T)$. Combining this with $|Q_{T}|\leq g(T)$, we get $\lim_{T\to\infty}\alpha(T)=0$. Also, since $|Q_{T}|\in o(f(T))$, there exists $T_{0}$ such that $|Q_{T}|\leq f(T)/2$ for all $T\geq T_{0}$. Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq 1$, we have
$\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ \>\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
@@
whenever $T\geq T_{0}$. Since $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})=1$ always, we have [^14] Note that Fubini’s theorem means we need not worry about the order of the expectation operators.
$\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu_{f,\mathbf{x}},\pi^{*})$
$=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})-\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\right]$
@@
$\lim_{T\to\infty}\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
@@
B.1 Proof roadmap

### v3

@@ 5.2 Proof sketch
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\sup\limits_{\pi\in\Pi}\min\limits_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\pi(s_{t})\neq\tilde{\pi}(s_{t}))\right]\in O\left(\sqrt{\frac{d\varepsilon T}{\sigma}\log T\log(1/\varepsilon)}+\frac{\varepsilon T}{\sigma}\log T\right)$
We will run a variant of Hedge on $\tilde{\Pi}$. The vanilla Hedge algorithm operates in the standard online learning model where on each time step, the agent selects a policy (or more generally, a hypothesis), and observes the loss of every policy. In general the loss function can depend arbitrarily on the time step, the policy, and prior events, but we will only use the indicator loss function $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$. Crucially, whenever we query and learn $\pi^{m}(s_{t})$, we can compute $\ell(t,\pi)$ for every $\pi\in\tilde{\Pi}$.
We cannot afford to query on every time step, however. Recently, [33] analyzed a variant of Hedge where losses are observed only in response to queries, which they call “label-efficient feedback”. They proved a regret bound when a query is issued on each time step with fixed probability $p$. Lemma 5.4 restates their result in a form that is more convenient for us (see Appendix B for details). Although their result is stated for non-adaptive adversaries, we explain in Section B.3 how their argument easily generalizes to adaptive adversaries. Full pseudocode for HedgeWithQueries can also be found in the appendix (Algorithm 2).
Lemma 5.4 (Lemma 3.5 in [33]).
Assume $\tilde{\Pi}$ is finite. Then for any loss function $\ell:[T]\times\tilde{\Pi}\to[0,1]$ and query probability $p$, HedgeWithQueries enjoys the regret bound
$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\pi)\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}$
where $\pi_{t}$ is the policy chosen at time $t$ and the expectation is over the randomness of the algorithm.
We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$ and combine this with Lemmas 5.1 and 5.3 (in the $\sigma$-smooth case) and with Lemma 5.2 (in the adversarial case). This yields a $O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$ bound on the number of mistakes made by Algorithm 1 (Lemma B.1).
The other key ingredient of the proof is analyzing the “ask for help when out-of-distribution” component. Combined with the local generalization assumption, this allows us to fairly easily bound the cost of a single mistake (Lemma B.2). The trickier part is bounding the number of resulting queries. It is tempting to claim that the states queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually not true. Instead, the packing analysis must be performed on a per-action basis, and we reuse Lemma B.1 to crucially show that there isn’t too much “spillover” between actions. The resulting overall query bound is given by Lemma B.7.
@@ 6 Conclusion and future work
In this paper, we proposed a model of avoiding catastrophe in online learning. We showed that achieving subconstant regret in our problem (with the help of a mentor and local generalization) is no harder than achieving sublinear regret in standard online learning.
There remain some technical questions within this paper’s model. One question is whether the time complexity of Algorithm 1 be improved, which currently stands at $\Omega(|\tilde{\Pi}|\cdot T)$ plus the time to compute the $\varepsilon$-cover. Also, we have not resolved whether our problem is tractable for finite VC dimension and fully adversarial inputs (although Appendix D shows that the problem is tractable for at least some classes with finite VC but infinite Littlestone dimension).
We are also interested in alternatives to the local generalization assumption, since an action which is safe in one state may not always be safe in a nearby state (depending on the definition of “nearby”). Some assumption is necessary, or there is no way to avoid costly mistakes without querying on every time step. One possibility is Bayesian inference. We intentionally avoided Bayesian approaches in this paper due to tractability concerns, but it seems premature to abandon those ideas entirely.
Finally, we are excited to apply the ideas in this paper to Markov Decision Processes (MDPs): specifically, MDPs where some actions are irreversible (“non-communicating”) and the agent only gets one attempt (“single-episode”). In such MDPs, the agent must not only avoid catastrophe but also obtain high reward. As discussed in Section 2, very little theory exists for RL in non-communicating single-episode MDPs. Can an agent learn near-optimal behavior in high-stakes environments while becoming self-sufficient over time? Formally, we pose the following open problem:
Is there an algorithm for non-communicating single-episode undiscounted MDPs which ensures that both the regret and the number of mentor queries are sublinear in $T$?
@@ Acknowledgements
This work was supported in part by a gift from Open Philanthropy to the Center for Human-Compatible AI (CHAI) at UC Berkeley. This paper also benefited from discussions with many other researchers. We would like to especially thank (in alphabetical order) Cameron Allen, Michael Cohen, Scott Emmons, Nika Haghtalab, Aly Lidayan, Juan Liévano-Karim, Bhaskar Mishra, and Matteo Russo.
@@ A.1 Proof roadmap
Recall that $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}$ denotes the expectation over both the random states and any additional randomness in the algorithm.
@@
    To complete the proof, we observe that $\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu_{f,\mathbf{x}},\pi^{m})$. This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\mathbf{x}}$ has high expected regret, then the worst case $\mu$ also has high expected regret.
@@
Lastly, while the intuition provided in Section 4.1 is accurate, the analysis will mostly occur in log space, so the bounds will look different. However, bounds of the form discussed in Section 4.1 can still be found as an intermediate step in Part 4 of the proof of Lemma A.4.
@@ A.2 Proof
Independently sample $\mathbf{x}\sim U(\{0,1\}^{f(T)})$ and $\mathbf{s}\sim U^{T}$. [^12] That is, the entire set $\{x_{1},\dots,x_{f(T)},s_{1},\dots,s_{T}\}$ is mutually independent. Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$,
@@
Part 3: $|V_{j}^{\prime}|\geq|V_{j}|/2$. Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in V_{j}$, so $a_{t}\in\{0,1\}$ for all $t\in V_{j}$. Since the agent chooses one of two actions for each $t\in V_{j}$, the more frequent action must be chosen chosen at least half of the time: $a_{t}=y_{j}$ for at least half of the time steps in $V_{j}$. Since $x_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $a_{t}=y_{j}\neq x_{j}$ for those time steps, so $|V_{j}^{\prime}|\geq|V_{j}|/2$.
@@
[…]
@@
Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$ for brevity. See 4.1
@@
If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{m}$. Let $f(T)=\sqrt{(g(T)+1)T}$. Then by Lemma A.1, $g(T)\in o(f(T))$ and $f(T)\in o(T)$. Combining this with $|Q_{T}|\leq g(T)$, we get $\lim_{T\to\infty}\alpha(T)=0$. Also, since $|Q_{T}|\in o(f(T))$, there exists $T_{0}$ such that $|Q_{T}|\leq f(T)/2$ for all $T\geq T_{0}$. Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq 1$, we have
$\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ \>\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
@@
whenever $T\geq T_{0}$. Since $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{m}(s_{t})=1$ always, we have [^13] Fubini’s theorem means we need not worry about the order of the expectation operators.
$\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu_{f,\mathbf{x}},\pi^{m})$
$=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{m}(s_{t})-\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\right]$
@@
$\lim_{T\to\infty}\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
@@
B.1 Context on Lemma 5.4
Before diving into the main proof, we provide some context on Lemma 5.4 from Section 5:
See 5.4
Lemma 5.4 is a restatement and simplification of Lemma 3.5 in [33]. First, [33] parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$. Second, [33] include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound. In our case, an expected query bound is sufficient, so we simply set $k=\hat{k}$. Third, [33] provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it. Fourth, their number of actions $n$ is equal to $|\tilde{\Pi}|$ in our setting. (Their actions correspond to policies in $\tilde{\Pi}$, not our actions in $A$.) Since [33] set $\eta=\max\Big{(}\frac{1}{T}\sqrt{\frac{\hat{k}\log n}{2}},\frac{k\hat{k}}{\sqrt{2}T^{2}}\Big{)}$, we end up with $\eta=\max\Big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\Big{)}$. Algorithm 2 provides precise pseudocode for the HedgeWithQueries algorithm to which Lemma 5.4 refers.
1:function HedgeWithQueries($p\in(0,1],\>\text{finite policy class }\tilde{\Pi},\>\text{unknown }\ell:[T]\times\tilde{\Pi}\to[0,1]$)
2:     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
3:     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
4:     for $t$ from $1$ to $T$ do
[…]
6:         if hedgeQuery then
7:              Query and observe $\ell(t,\pi)$ for all $\pi\in\tilde{\Pi}$
8:              $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
9:              $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
10:              Select policy $\arg\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
11:         else
12:              $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
13:              Sample $\pi_{t}\sim P$
14:              Select policy $\pi_{t}$
Algorithm 2 A variant of the Hedge algorithm which only observes losses in response to queries.
B.2 Main proof